OpenStudy (anonymous):

Solve the system of equations. 4x+3y=13 y=-x+4

OpenStudy (anonymous):

@Whiterabbit541 @ShizukaTheOtaku @hwyl

OpenStudy (anonymous):

@Nurali @Abhisar

OpenStudy (anonymous):

@Projects

OpenStudy (projects):

I will try to solve your system of equations. c: y=−x+4;4x+3y=13 Step: Solve y=−x+4 for y: Step: Substitute (−x+4) for y in 4x+3y=13: 4x+3y=13 4x+3(−x+4)=13 x+12=13(Simplify both sides of the equation) x+12+−12=13+−12(Add -12 to both sides) x=1 Step: Substitute (1) for x in y=−x+4: y=−x+4 y=−1+4 y=3(Simplify both sides of the equation) Your answer, y=3 and x=1